OpenStudy (anonymous):

What are all the real zeros of y = (x - 12)3 - 10?

OpenStudy (anonymous):

@vineeth10

OpenStudy (anonymous):

any options?

OpenStudy (anonymous):

yes hold up

OpenStudy (anonymous):

I looked it up and "real zeros" is where the graph touches the x-axis, meaning y = 0. So to solve this you set y = 0, and then solve for x. I already solved it and got the answer, but try it yourself and see what you get.

OpenStudy (anonymous):

y = (x-12)3 - 10 ... y = 3x - 36 - 10 y = 3x - 46 ... 3x - 46 = 0

OpenStudy (anonymous):

3x = 46 x = 46/3 or 15 1/3

OpenStudy (anonymous):

Yes the correct answer is x = 46/3. vineeth10 is correct =)

OpenStudy (anonymous):

yep thanks @lamaguineapig

OpenStudy (anonymous):

B.x=^3sqrt 10 +12

OpenStudy (anonymous):

C.x=^3sqrt-10+12

OpenStudy (anonymous):

D.x=^3sqrt 12-10

OpenStudy (anonymous):

OH u meany = (x - 12)^3 - 10?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i mean mean sorry

OpenStudy (anonymous):

(x−12)^3−10=0 (x−12)^3=10 (x−12)=^3√10 x=^3√(10)12

OpenStudy (anonymous):

x=^3√(10)+12

OpenStudy (campbell_st):

why not make it easy set the equation to zero \[(x - 12)^3 - 10 = 0\] so \[(x - 12)^3 = 10\] then \[x - 12 = \sqrt[3]{10}\] and you can finish from here

OpenStudy (anonymous):

i am right? @campbell_st

OpenStudy (campbell_st):

looks like it to me...

OpenStudy (anonymous):

B is the answer